OpenStudy (anonymous):

Solve for x: |x − 2| + 10 = 12 x = 0 and x = 4 x = −4 and x = 0 x = −20 and x = 4 No Solution

OpenStudy (anonymous):

Thanks So Much For Your Help!

OpenStudy (lgbasallote):

i'll try giving an example again... |x+3| - 2 = 4 first step is to add 2 |x+3| = 4+2 |x+3| = 6 now i solve for x when x + 3 = 6 and x+3 = -6 when x + 3 = 6 x = 6 - 3 x = -3 when x + 3 = -6 x = -6 - 3 x = -9 does that help?

OpenStudy (anonymous):

|x − 2| + 10 = 12 means x-2+10=12 or 2-x+10=12 x= 12-10+2 x=4 or -x = 12-10-2 -x = 0 x = 0

OpenStudy (anonymous):

@gitahimart Can You Help Me A Little More?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

I Will Mention You When I Need Your Help.